### v2

# Spatial Reasoning with Vision-Language Models in Ego-Centric Multi-View Scenes

Abstract
Understanding 3D spatial relationships remains a major limitation of current Vision-Language Models (VLMs). Prior work has addressed this issue by creating spatial question-answering (QA) datasets based on single images or indoor videos. However, real-world embodied AI agents—such as robots and self-driving cars—typically rely on ego-centric, multi-view observations. To this end, we introduce Ego3D-Bench, a new benchmark designed to evaluate the spatial reasoning abilities of VLMs using ego-centric, multi-view outdoor data. Ego3D-Bench comprises over 8,600 QA pairs, created with significant involvement from human annotators to ensure quality and diversity. We benchmark 16 SOTA VLMs, including GPT-4o, Gemini1.5-Pro, InternVL3, and Qwen2.5-VL. Our results reveal a notable performance gap between human level scores and VLM performance, highlighting that current VLMs still fall short of human level spatial understanding. To bridge this gap, we propose Ego3D-VLM, a post-training framework that enhances 3D spatial reasoning of VLMs. Ego3D-VLM generates cognitive map based on estimated global 3D coordinates, resulting in 12% and 56% average improvements on multi-choice QA and absolute distance estimation, respectively. Ego3D-VLM can be integrated with any existing VLM. Together, Ego3D-Bench and Ego3D-VLM offer valuable tools for advancing toward human level spatial understanding in real-world, multi-view environments.
Figure 1: Ego3D-Bench is a 3D spatial benchmark for VLMs using ego-centric multi-view images. It spans ego- and object-centric perspectives across 5 categories. A significant gap exists between human and VLM performance; our method, Ego3D-VLM, consistently narrows this gap.
1 Introduction
3D spatial understanding is a critical capability for embodied AI agents operating in the real world [4, 40]. This includes perceiving the location of surrounding objects, estimating their distances, and reasoning about their motion. VLMs have recently emerged as powerful tools to integrate visual perception and language reasoning, making them promising components for building intelligent embodied AI systems [28, 23]. Thus, enhancing and evaluating the spatial understanding of VLMs has become an increasingly important research direction [38, 6, 5, 51, 50, 39, 42, 41, 36, 33].
Recent benchmarks on spatial understanding focus mainly on spatial reasoning from single images [6, 5, 24] or videos captured in static indoor environments [44, 8]. In such cases, spatial understanding is framed around passive observations, where a single camera moves in a room to create a video, and the model should infer spatial relationships in a relatively static scene. This setup differs fundamentally from the perceptual experience of real-world embodied agents such as self-driving cars or mobile robots. These agents rely on ego-centric multi-view observations [28], provided by multiple cameras simultaneously capturing front, side, and rear views of their surroundings. These views are not interchangeable or purely visual; they carry explicit spatial semantics tied to the agent’s frame of reference. E.g, “left” & “right” refer to fixed directions relative to the agent’s body, and must be interpreted consistently over time as the agent moves through dynamic environments. This distinction is crucial: while prior video-based datasets may offer multiple viewpoints, they do not reflect the structured, directional, and temporally evolving nature of ego-centric multi-view inputs. Also, existing benchmarks do not evaluate VLMs’ reasoning ability across these spatially grounded perspectives in dynamic, real-world scenes.
This gap motivates the need for new benchmarks that better align with the spatial reasoning demands of embodied agents. To this end, we introduce Ego3D-Bench, a benchmark of 8.6K QA pairs carefully curated from the validation set of 3 public datasets: NuScenes [2], Waymo Open Dataset [37], and Argoverse 1 [3]. Human annotators played a central role in both the dataset construction and the rigorous quality review process to ensure the reliability of the benchmark. We focused specifically on ego-centric multi-view tasks, rather than building a general-purpose benchmark, to complement existing monocular spatial benchmarks (e.g., VSI-Bench [44]). Thus, we excluded questions that can be answered based on each view independently across multiple images (e.g., counting objects in each image) or general knowledge of LLMs (e.g., estimating the size of well-known objects). To our knowledge, Ego3D-Bench is the first benchmark to evaluate spatial reasoning of VLMs given ego-centric multi-view inputs.
We evaluated 16 SOTA VLMs, including generalist and 3D spatial ones on Ego3D-Bench, revealing a significant gap between human performance and current VLMs. We hypothesize that a key limitation lies in the inability of VLMs to construct a coherent world model from multi-view images. In contrast, humans naturally integrate visual info from their left, right, and front views into a unified spatial representation, enabling real-time reasoning and navigation. Prior work has attempted to bridge this gap by first generating a 3D point-cloud [38, 19, 18, 47, 17] or rendering a bird-eye-view (BEV) image of the scene [31]. Although point-clouds and BEV images offer rich spatial information, they are challenging to reconstruct in dynamic environments, struggle with sparse multi-view inputs, and significantly increase inference time—often by a factor of ten [53].
To this end, we propose, Ego3D-VLM, a post-training method that improves 3D spatial understanding of VLMs. The main idea of Ego3D-VLM is to create a textual cognitive map of the surrounding. The textual cognitive map defines a coordinate system center on the ego and locates important object (i.e., those referred to in the input prompt) in 3D coordinate space. Unlike point-clouds and BEV image methods [38, 31], our cognitive map only focuses on referred objects, making the number of input tokens significantly smaller and enabling efficient reasoning.
Given multi-view images as input, we first use referring expression comprehension (REC) models to find the 2D location of referred expressions in pixel space. We also use a metric depth estimator to estimate the depth values. We then convert the 2D points to 3D points in camera coordinate space and transform 3D points from all views to the global coordinate space (i.e., front camera coordinate space). We define a cognitive map generator function that returns a textual cognitive map given 3D coordinates. The textual cognitive map defines a coordinate space for the VLM and organizes detected objects (expressions) based on the view-point. The textual cognitive map consistently improves SOTA VLMs on spatial reasoning. The major contributions of our work are:
  • We introduce Ego3D-Bench, an ego-centric multi-view benchmark to evaluate 3D spatial understanding of VLMs.
  • We propose Ego3D-VLM, a plug-and-play post-training framework to enhance 3D spatial understanding of VLMs, especially in ego-centric multi-view scenarios.
  • Through extensive experiments, we demonstrate that Ego3D-VLM significantly improves SOTA generalist as well as 3D spatial VLMs in 3D reasoning.
2 Related Work
3D Spatial Benchmarks/Datasets for VLMs. We categorize prior datasets and benchmarks into three groups based on the input data type. (1) [25, 20, 24, 36, 50, 6, 10] focus on single-view images. (2) [44, 8, 22, 45, 33] focus on videos captured from indoor static scenes. As noted, this setup differs from the perceptual experience of real-world embodied agents, which involves ego-centric multi-view images. (3) All-Angle [46] is the first multi-view benchmark for VLMs. However, in their setup, multiple cameras observe a scene from different directions—similar to surveillance or motion capture systems. In contrast, Ego3D-Bench is the first ego-centric multi-view benchmark for VLMs created from dynamic outdoor scenes.
Figure 2: Overview of our Ego3D-Bench creation pipeline. Human annotators played a key role in the process. Right: the distribution of the samples. Ego.: Ego-centric, Obj.: Object-centric.
3D Spatial VLMs. These models—also referred to as 3D-LLMs or 3D-MLLMs—aim to perform tasks such as 3D grounding, spatial reasoning, depth estimation, and distance measurement. We categorize existing approaches into two main groups: (1) models that take point-clouds as input or reconstruct them from multi-view images, and (2) models that operate directly on image data.
[18, 17, 19, 38, 49, 47] fall in the first group. While point-cloud representations offer rich spatial information, they are difficult to reconstruct in dynamic scenes, often struggle with sparse multi-view data, and significantly increase inference time by over 10×. The 2nd group includes LLaVA-3D [53], Video-3D LLM [52], GPT4Scene [31], MM-Spatial [8], SpatialVLM [5], SpatialRGPT [6], and SpatialPIN [27]. Our work falls in this image-based category.
LLaVA-3D and Video-3D LLM use depth maps and camera poses for 3D positional encoding. GPT4Scene leverages BEV to address 3D queries. However, these models are primarily trained on indoor, static scenes and limited in handling quantitative spatial relationships such as object distances. SpatialVLM addresses this limitation by introducing a synthetic data generation pipeline, enabling more robust spatial reasoning [34]. SpatialRGPT enhances input representation using region proposals alongside the original image. MM-Spatial proposes a VLM that supports Chain-of-Thought spatial reasoning involving 2D grounding and depth estimation, and can also leverage depth input via tool-use. Different from SpatialRGPT, SpatialVLM, and MM-Spatial, our proposed Ego3D-VLM is a post-training method, can be applied to any existing VLM, and enhances spatial understanding of VLMs and outperforms prior works on ego-centric multi-view reasoning.
3 Ego3D-Bench
Ego3D-Bench is designed to quantitatively evaluate 3D spatial understanding of VLMs from multi-view outdoor images. A major distinction between Ego3D-Bench and previous works is its focus on ego-centric multi-view data—an approach particularly relevant for applications in self-driving and robotics. Ego3D-Bench contains over 8.6K QA pairs in 5 distinct categories. This benchmark is constructed from the validation sets of 3 prominent outdoor multi-view datasets: nuScenes [2], Waymo Open Dataset [37], and Argoverse 1 [3], featuring 6, 7, and 5 distinct camera views, respectively (Figure 2). These datasets cover diverse outdoor environments, including urban areas, highways, and rural regions. Ego3D-Bench leverages the multi-view nature of these datasets to formulate questions that require fine-grained visual comprehension across different viewpoints as well as reasoning over 3D spatial relationships. In this section, we describe the detailed process of constructing the questions of Ego3D-Bench.
Figure 3: Ego- and object-centric samples from each category in Ego3D-Bench.
3.1 Benchmark Construction
Creation of Source Files: Outdoor datasets, unlike their indoor counterparts, often feature multiple instances of the same object within each scene. While indoor environments typically contain unique items—such as a single oven or television per room—outdoor scenes commonly include numerous similar objects, like multiple cars or pedestrians. This makes it challenging to uniquely reference a target object in the scene. Thus, annotators begin by carefully reviewing each scene to identify unique objects (called "Filtering" in Figure 2). They then compose concise captions describing these objects (called "Captioning" in Figure 2). The descriptions are designed to be short, yet discriminative such that each object can be uniquely identified. Furthermore, the ground-truth 3D annotations (bounding boxes) are collected from the source datasets. Object captions, 3D annotations, and the camera view from which the object is visible are used to construct a source file for QA creation.
Creation of QAs: Each question category follows a predefined template with placeholders for object names, such as How far is <obj1>in <view1>from <obj2>in <view2>? (see Appendix A.2 for all question templates). Each question is constructed by placing the generated object annotation and camera views from the source file in the question template. To generate the answers, we use rule-based functions that leverage ground-truth 3D annotations. For challenging categories, such as motion reasoning, a final human review is conducted to ensure the accuracy of the QA pairs.
3.2 Benchmark Details
Figure 3 illustrates different question categories in Ego3D-Bench. We emphasize on questions that require understanding the 3D world by utilizing information from multiple views. Thus, we exclude questions that can be answered using a single-view or by analyzing each view independently (e.g., counting the number of objects), as they do not contribute to multi-view spatial reasoning. We define question from the perspective of both ego and objects in the scene. To clearly indicate the perspective of each question, we categorize them into ego-centric or object-centric. In the following, we describe the five tasks (each composed of ego-centric and object-centric types) in our benchmark.
(1) Absolute Distance Measurement. This category asks the VLM to estimate the absolute metric distance between the ego and another object in the scene or between two different objects from different views. This category is designed in two forms of multi-choice QA and absolute meter.
(2) Relative Distance Measurement. In this task, the VLM is asked to determine which of two objects is closer—either to the ego or to a specific object, designed in the form of two-choice QA.
(3) Localization. This category is only object-centric and assesses the VLM’s ability to localize objects within a scene. Specifically, the model is asked to infer the location of object-1 from the perspective of object-2 in the form of multi-choice QA.
(4) Motion Reasoning. This category defines a coordinate system using cardinal directions and asks the VLM that if the ego or object-1 moves in a direction, whether or not it gets closer to or farther away from object-2. Answering this yes/no question requires mapping the spatial relationship between the objects and how the distances would change if one object moves in a specific direction.
(5) Travel Time. Given a specific motion speed, this category asks (via multi-choice QA) to estimate the required time to move from the location of either the ego or object-1 to the location of object-2.
Evaluation Metrics. We design most of the questions as multi-choice QA and use accuracy as the evaluation metric. We also have two absolute distance estimation for which we use root mean squared error (RMSE) in meters as the evaluation metric.
Figure 4: An overview of Ego3D-VLM, our post-training 3D spatial understanding framework.
4 Post-Training 3D spatial understanding : Ego3D-VLM
Figure 4 shows an overview of our proposed framework, Ego3D-VLM. Given a set of multi-view images and a natural language prompt, we use a REC model to detect all objects mentioned in the prompt. For each camera view $v\in\{1,2,\dots,V\}$, the REC model returns a set of detected objects:
$\mathcal{O}^{(v)}=\left\{\left(b_{i}^{(v)},c_{i}^{(v)}\right)\right\}_{i=1}^{N^{(v)}},$ (1)
where $b_{i}^{(v)}\in\mathbb{R}^{4}$ denotes the 2D bounding box coordinates for the $i$-th object in view $v$, and $c_{i}^{(v)}$ is the corresponding referring expression match. We compute the 2D pixel-space center of each bounding box as $\mathbf{u}_{i}^{(v)}\in\mathbb{R}^{2}$ to create a list of 2D centers of the objects in the pixel space.
Camera Coordinate Transformation. To obtain 3D spatial info, we use a metric depth estimator to predict a dense depth map $D^{(v)}\in\mathbb{R}^{H\times W}$ for each view $v$. For each detected object center $\mathbf{u}_{i}^{(v)}=(x_{i},y_{i})^{\top}$, we extract the corresponding depth value $d_{i}^{(v)}=D^{(v)}(x_{i},y_{i})$. We then project each center point into the 3D camera coordinate system using the camera intrinsics $K^{(v)}\in\mathbb{R}^{3\times 3}$:
$\mathbf{p}_{\text{cam},i}^{(v)}=d_{i}^{(v)}\cdot\left(K^{(v)}\right)^{-1}\begin{bmatrix}x_{i}\\ y_{i}\\ 1\end{bmatrix}\in\mathbb{R}^{3}.$ (2)
Next, we transform the 3D point from the local to the global coordinate system, i.e., defined as the front camera view point coordinate system. This replicates the human perception mechanism given multi-view cameras by using the front view as the reference building the 3D world based on that. Using the rotation matrix $R^{(v)}\in\mathbb{R}^{3\times 3}$ and translation vector $T^{(v)}\in\mathbb{R}^{3\times 3}$, we have:
$\mathbf{p}_{\text{global},i}^{(v)}=\begin{bmatrix}R^{(v)}&T^{(v)}\\ 0&1\end{bmatrix}\cdot\begin{bmatrix}\mathbf{p}_{\text{cam},i}^{(v)}\\ 1\end{bmatrix}\in\mathbb{R}^{4}.$ (3)
This process effectively simulates human spatial perception by leveraging multi-view images to construct a unified 3D representation of the scene.
Relational Scaling. Humans estimate object sizes using known references—e.g., knowing a person is $\approx$ 1.7 m tall helps infer the size of nearby objects. Inspired by this, we scale 3D points $\mathbf{p}_{\text{global},i}^{(v)}$ using familiar object categories (e.g., sedans, humans, bikes) identified in a few representative frames across all camera views. We compute the average observed height $h_{\text{est}}$ and scale all 3D points by $s={h_{\text{cs}}}/{h_{\text{est}}}$, where $h_{\text{cs}}$ is the canonical common sense height (e.g., 1.7 m for humans). This yields $\mathbf{p}_{\text{scaled},i}^{(v)}=s\cdot\mathbf{p}_{\text{global},i}^{(v)}$, producing physically plausible scales without ground-truth depth.
Creating a Cognitive Map. We define a cognitive map generator function $\mathcal{F}_{\text{cog}}$, which takes as input the set of 3D global coordinates and corresponding referring expressions for all detected objects across all camera views. Specifically, for each object $i$ detected from view $v$, we denote its global 3D position as $\mathbf{p}_{\text{global},i}^{(v)}$ and its matched referring expression as $c_{i}^{(v)}$. The function outputs a textual cognitive map $C$, defined as:
$C=\mathcal{F}_{\text{cog}}\left(\left\{\left(\mathbf{p}_{\text{global},i}^{(v)},c_{i}^{(v)}\right)\right\}_{i,v}\right).$ (4)
$\mathcal{F}_{\text{cog}}$ constructs an ego-centric world model centered on the agent. It integrates multi-view detections and links each referred object to its spatial position and originating viewpoint. The resulting cognitive map captures both linguistic references and spatial semantics in a compact, human-interpretable form—enabling grounded reasoning and situational awareness. Figure 5 illustrates a sample prompt, the corresponding multi-view images, and the generated cognitive maps.
Figure 5: Example question with associated textual and visual cognitive maps.
Given a VLM $\mathcal{V}$ that answers queries by visual-textual contexts, it takes as input the cognitive map $C$, a set of multi-view images $\mathcal{I}=\{I^{(v)}\}_{v}$, and a natural language query $q$, to return an answer $a$:
$a=\mathcal{V}(C,\mathcal{I},q).$ (5)
The cognitive map $C$ provides structured spatial grounding, while $\mathcal{I}$ supplies rich visual cues—such as appearance, color, and fine-grained context—not captured in $C$. Together, they guide the VLM in interpreting and answering the query.
5 Evaluation on Ego3D-Bench
In this section, we present a comprehensive evaluation of VLMs on Ego3D-Bench. We organize our experiments into 4 parts by benchmarking (1) generalist VLMs (i.e., models trained for general vision-language tasks), (2) 3D-VLMs (i.e., models trained specifically for 3D SU), (3) VLM+Depth+REC (i.e., generalist VLMs augmented with depth and REC tools), and (4) ablation studies. We use a fixed R1-style prompt [16] in the evaluations of all models (details in the Appendix A.8). We use Grounding-Dino-Base [35] as the REC model and Depth-Anything-V2-Metric-Large [7] as the metric depth estimator in all experiments.
Table 1: Comparison results of generalist VLMs vs. Ego3D-VLM on Ego3D-Bench.
[…]
Ego
Obj.
Ego
Obj.
Travel
Ego
Obj.
Ego
Obj.
Model
Dist.
Dist.
Loc.
Mot.
Mot.
Time
Rel.
Rel.
[…]
Dist.
Dist.
[…]
Human Level 78.2 57.1 100 100 95.3 72.7 85.0 94.1 85.3 - - -
Chance Level 25.8 25.5 24.3 49.7 50.6 24.0 49.9 50.1 37.5 - - -
Closed-source Models
GPT-4o 51.4 38.4 47.7 86.9 71.1 35.8 65.9 84.2 56.7 8.5 29.5 19.2
Ego3D-VLM GPT-4o 76.3 58.9 57.3 87.3 89.7 59.9 70.6 85.3 73.2 6.3 8.4 7.4
Gemini-1.5-Pro 51.3 35.8 49.4 88.4 70.6 37.7 55.5 71.4 57.5 10.7 28.6 19.6
Ego3D-VLM Gemini-1.5-Pro 65.0 62.0 67.2 93.9 91.4 58.4 66.2 80.6 73.1 6.6 7.8 7.2
Qwen2.5 Family
Qwen2.5-3B 21.5 29.4 28.8 50.3 41.9 30.9 54.1 56.1 39.1 30.5 33.7 32.1
Ego3D-VLM Qwen2.5-3B 35.0 30.0 29.8 52.4 56.4 29.6 60.1 62.0 44.4 12.7 12.5 12.6
Qwen2.5-7B 32.7 31.5 30.5 45.9 44.0 34.5 43.2 66.5 41.1 25.1 35.5 30.3
Ego3D-VLM Qwen2.5-7B 59.4 54.5 33.1 62.3 58.2 49.4 50.5 66.9 54.3 8.1 10.9 9.5
Qwen2.5-32B 45.4 40.7 49.6 75.6 74.1 40.1 54.0 79.0 57.3 21.2 10.4 15.8
Ego3D-VLM Qwen2.5-32B 63.7 62.6 54.5 87.7 86.2 40.8 62.6 72.0 65.5 11.6 15.9 13.7
Qwen2.5-72B 42.4 38.6 54.8 86.8 68.9 38.5 53.3 80.5 58.0 10.3 22.2 16.2
Ego3D-VLM Qwen2.5-72B 62.1 61.4 58.2 94.0 84.5 56.0 63.1 76.6 69.5 6.8 8.3 7.5
InternVL3 Family
InternVL3-8B 25.8 28.7 29.8 54.1 54.8 36.1 49.9 65.2 43.1 15.2 39.1 27.2
Ego3D-VLM InternVL3-8B 65.4 56.1 37.0 73.0 71.4 49.0 63.5 66.0 60.1 6.8 9.0 8.0
InternVL3-14B 46.0 35.6 35.9 63.2 65.9 41.6 55.5 70.1 51.7 10.6 24.5 17.6
Ego3D-VLM InternVL3-14B 70.3 60.9 50.5 79.8 83.1 50.2 63.0 70.7 66.1 6.6 8.8 7.7
InternVL3-38B 35.4 31.0 39.4 66.6 64.9 38.0 61.0 77.3 51.7 8.6 42.2 25.4
Ego3D-VLM InternVL3-38B 55.1 64.5 53.4 87.2 88.9 51.9 66.6 76.5 68.0 8.2 8.7 8.5
InternVL3-78B 54.6 48.4 50.3 77.7 70.0 44.8 57.0 76.6 59.9 12.0 15.5 13.8
Ego3D-VLM InternVL3-78B 68.3 62.7 62.9 91.6 89.2 55.1 66.3 78.2 71.8 6.8 8.1 7.4
Ovis2 Family
Ovis2-4B 29.8 28.9 18.4 47.5 48.1 36.9 54.2 70.3 41.8 17.1 29.5 23.3
Ego3D-VLM Ovis2-4B 62.1 46.9 20.1 49.1 51.9 33.3 62.0 60.4 48.2 6.5 10.4 8.5
Ovis2-8B 25.6 28.6 31.1 45.3 51.4 31.4 50.7 68.2 41.5 11.5 30.2 20.8
Ego3D-VLM Ovis2-8B 64.9 54.9 33.5 57.2 57.1 46.1 65.1 71.0 56.2 6.0 9.5 7.8
Ovis2-16B 41.7 36.5 41.1 50.5 52.5 27.9 50.9 74.9 47.0 10.8 16.6 13.7
Ego3D-VLM Ovis2-16B 63.4 58.1 43.7 53.9 73.4 51.1 61.0 82.9 60.9 6.6 8.3 7.4
Chance/Human Performance Levels. We provide frequency-based random selection as the chance level baseline for the multi-choice questions. Furthermore, we conduct human evaluation on multi-choice questions, where 10% of the questions from each category are randomly sampled and evaluated by each human annotator. Table 1 presents the results of this analysis. While humans can accurately answer the questions that require reasoning about relative location of the objects in space, their performance degrades in questions that require estimation of the exact distance between objects. This highlights the challenging nature of accurate distance estimation.
5.1 Benchmarking Generalist VLMs
We use GPT-4o [30] and Gemini-2-Flash [9] closed-source VLMs, and 3 competitive families of open-source models: InternVL3 [11], Qwen2.5-VL [32], and Ovis2 [26] open-source models. Inference time analysis, numerical results on more generalist VLMs, and qualitative examples are given in the Appendix A.5, A.4, and A.8.
As given in Table 1, the performance of the VLMs varies considerably across different model parameter sizes. Smaller models (e.g., 3B and 8B) operate near chance level, indicating limited capacity for multi-view 3D reasoning. In contrast, larger models demonstrate substantial improvements over the chance level, yet still exhibit a noticeable gap when compared to human level performance. Furthermore, our proposed Ego3D-VLM provides significant improvement across all model sizes and tasks (average 56% relative improvement in RMSE and 12% in Accuracy), underscoring the importance of providing structured spatial representations to the model in 3D understanding tasks.
Figure 6: Average performance of leading VLMs w/ and w/o Ego3D-VLM vs. chance & human levels on each category of Ego3D-Bench.
Performance Analysis. Figure 6 shows the average performance of 4 leading models (GPT-4o, Gemini-1.5-Pro, InternVL3-78B, and Qwen2.5-VL-72B) with and without Ego3D-VLM integration across all multiple-choice QAs. Travel time, localization, and object-centric absolute distance are the most challenging tasks for VLMs with 40%-45% average accuracy. We argue that these categories require intricate spatial reasoning and the ability to construct an accurate mental map by the relative positioning of objects. While humans can simply build such maps and achieve perfect accuracy, VLMs struggle to replicate this level of SU, indicating a key area for further development. In the absolute distance tasks, the incorporation of 3D location data through the cognitive map substantially narrows the gap between model and human level accuracy. Notably, for object-centric absolute distance, VLMs augmented with Ego3D-VLM even surpass human. This is expected, as human estimation of 3D distances in object-centric cases is prone to substantial error without explicit 3D spatial info. Conversely, in the localization task, VLMs continue to fall short of human proficiency, even with cognitive map support.
Table 2: Blind Ego3D-VLM performance analysis.
Only Mult. Choice Abs. Ans.
[…]
InternVL3-8B 43.1 27.2
Ego3D-VLMInternVL3-8B 53.1 11.4
Ego3D-VLMInternVL3-8B 60.1 8.0
InternVL3-14B 51.7 17.6
Ego3D-VLMInternVL3-14B 55.6 10.8
Ego3D-VLMInternVL3-14B 68.0 8.5
Blind VLM Performance. This baseline evaluates how much spatial reasoning can be achieved by VLMs using only textual input, without any visual information, relying solely on their world knowledge [29]. We report the average performance of GPT-4o and Gemini-1.5-Pro models. The blind VLMs perform 5% worse than vision-enabled VLMs (53.8% vs. 58.8%) and 16.4% better than chance level. These results are consistent with findings from prior single-view and video-based spatial reasoning benchmarks [6, 44].
Blind Ego3D-VLM Performance. we use the same models for fair comparison. Table 2 shows that Ego3D-VLM with VLM outperforms Ego3D-VLM with LLM since VLMs can ignore false positives in the cognitive map and remain robust when false negatives occur, whereas LLMs suffer performance drops in those cases.
5.2 Benchmarking 3D-VLMs
3D-VLM models have been trained on datasets designed for 3D SU, such as absolute distance estimation and relative location inference. We benchmark SpatialRGPT and two checkpoints trained with the Spatial-VLM framework [5]: SpaceThinker-Qwen2.5-3B [34] and SpaceQwen2.5-3B. The SpatialRGPT [6] model assumes that specific regions of the image are annotated with bounding boxes and is trained to answer 3D questions based on these regions. To evaluate SpatialRGPT on our Ego3D-Bench, we reformat the input: object names in the questions are replaced with placeholder labels (e.g., region-i), and a list of corresponding region captions is passed to the REC model to generate bounding boxes. These estimated bounding boxes were then overlaid on the images before being fed to SpatialRGPT.
Table 3 presents the performance of the above-mentioned 3D-VLMs on our Ego3D-Bench. SpaceThinker-Qwen2.5-3B achieves the highest overall performance, surpassing both SpatialRGPT with 8B parameters and other generalist VLMs of similar scale. This outcome highlights the critical role of dedicated 3D spatial pretraining and end-to-end architecture in advancing VLM capabilities for spatial reasoning tasks. Moreover, augmenting SpaceThinker with Ego3D-VLM leads to an average improvement of 3% on multiple-choice questions and a reduction of more than 4 meters in absolute distance RMSE, emphasizing the effectiveness of our proposed solution.
Table 3: Comparison results of 3D-VLMs and our method (Ego3D-VLM) on Ego3D-Bench.
[…]
Ego
Obj.
Ego
Obj.
Travel
Ego
Obj.
Ego
Obj.
Model
Dist.
Dist.
Loc.
Mot.
Mot.
Time
Rel.
Rel.
[…]
Dist.
Dist.
[…]
SpaceQwen2.5-VL-3B 18.1 21.4 16.1 35.6 35.2 30.2 31.2 32.0 27.5 10.6 15.7 13.2
SRGPT-VILA1.5-8B 45.7 35.5 39.6 43.6 45.8 24.9 50.8 71.7 44.7 11.5 15.1 13.3
SThinker-Qwen2.5-3B 38.9 39.0 21.9 57.5 53.4 27.1 52.8 71.6 45.2 15.2 16.9 16.0
Ego3D-VLMSThinker-Qwen2.5-3B 50.6 44.2 26.4 53.4 57.0 42.6 54.7 59.8 48.6 11.1 12.2 11.6
5.3 Benchmarking VLM+Depth+REC.
Table 4: Ego3D-VLM vs. generalist VLMs with REC+depth tools on Ego3D-Bench.
Mult. Choice Abs. Ans.
[…]
InternVL3-8B 43.1 27.2
InternVL3-8B+Depth+REC 51.6 13.1
Ego3D-VLM InternVL3-8B 60.1 8.0
Qwen2.5-7B 41.1 30.3
Qwen2.5-7B+Depth+REC 49.4 11.8
Ego3D-VLM Qwen2.5-7B 54.3 9.5
This category enhances the base VLM with a metric depth estimator and a REC model. We pass each query to the REC model [35] to extract BBox of the referred objects, and use a metric depth estimator [7] to estimate depth values. We then construct a list that pairs each matched referring expression with its corresponding BBox and estimated depth (i.e., distance from the ego). An example entry in this list is: [Front-View: Detected pedestrian with red hat at bbox [x1, y1, x2, y2], depth: z, Back-View:...] (examples in the Appendix A.8). As shown in Table 4, equipping the VLM with REC and depth estimator models improves its baseline performance, highlighting the base model’s limitations in accurately identifying objects and estimating their depths. However, even with these enhancements, it falls short compared to our proposed Ego3D-VLM framework. This underscores the importance of integrating depth information into a unified map representation, enabling the VLM to reason more effectively about the 3D relationships between objects across different views.
5.4 Ego3D-VLM on Further Benchmarks
Our Ego3D-VLM is designed for ego-centric multi-view cases, which are particularly relevant for AI agent applications. In this section, however, we evaluate the performance of Ego3D-VLM in alternative multi-view settings, specifically those used in All-Angle Bench [46] and VSI-Bench [44]. Table 5 presents the performance of Ego3D-VLM compared to open- and closed-source baseline VLMs on All-Angle Bench and VSI-Bench. Despite the differences in data settings compared to our primary benchmark, Ego3D-VLM still outperforms the respective baselines, demonstrating its adaptability across diverse multi-view scenarios. More details are given in the Appendix A.5.
Table 5: Comparison results on All-Angle Bench and VSI-Bench (accuracy).
Model All-Angle-Bench VSI-Bench
GPT-4o 47.8 34.0
Gemini-1.5-Pro 47.4 45.4
Gemini-1.5-Flash 46.6 42.1
InternVL3-8B 47.9 38.1
Ego3D-VLMInternVL3-8B 49.5 39.6
InternVL3-14B 50.3 38.2
Ego3D-VLMInternVL3-14B 52.1 40.0
Table 6: Ablation on Ego3D-VLM main components. v4: all components.
Mult. Choice Abs. Ans.
[…]
v0 InternVL3-8B 43.1 27.2
v1 + CogMap (est. $R,T,K$) 56.0 10.8
v2 + $K$ 56.3 10.1
v3 + $R,T$ 58.4 10.4
v4 + Scaling 60.1 8.0
v5 + List of objects 61.8 6.5
v6 GT CogMap 79.4 1.3
Table 7: Perception-reasoning disentanglement.
GT Mult. Choice Abs. Ans.
[…]
InternVL3-8B 43.1 27.2
InternVL3-8B 50.2 21.0
Ego3D-VLMInternVL3-8B 60.1 8.0
Ego3D-VLMInternVL3-8B 62.2 6.8
InternVL3-14B 51.7 17.6
InternVL3-14B 51.8 16.2
Ego3D-VLMInternVL3-14B 66.5 7.7
Ego3D-VLMInternVL3-14B 66.8 7.5
5.5 Ablation Studies
Ego3D-VLM Components. Table 6 presents an ablation on Ego3D-VLM core components. Starting from the baseline, v0, we incrementally add each component. In v1, we use a cognitive map with estimated rotation ($R$), translation ($T$), and intrinsic params ($K$). Specifically, all cameras are positioned at the coordinate center with $T=[0,0,0]$. The front camera uses an identity rotation matrix, while all the other cameras—front-right, right, back-right, back, etc. —are incrementally rotated 45° around the Y-axis. The focal length is estimated from the images’ approximate field of view. In v2, we use the actual $K$ and in v3 we further use actual $R$, and $T$ relative to the front camera, i.e., often available as a fixed parameter for an embodied AI agent. v3 and v1 obtain a comparable RMSE, while v3 is only 2.4% higher in multi-choice QAs. Thus, even with estimated camera parameters our cognitive map can significantly boost the baseline’s performance. v4 adds relational scaling to cognitive map which decreases the RMSE by 2.5 meters. v4 is indeed Ego3D-VLM with all components. We provide two more ablations to evaluate the upper-bound of Ego3D-VLM. v5 assumes that the list of objects (only their names) in the input query are provided which enhances the REC results. v6 is the ground-truth cognitive map with ground-truth 3D locations.The difference between v6 and human level is only 5% showing the upper bound of Ego3D-VLM.
Table 8: Ablation on cognitive map format.
Mult. Choice Abs. Ans.
[…]
Visual Cog-Map 50.9 14.4
JSON Cog-Map 60.0 8.4
Textual Cog-Map 60.1 8.0
Perception-Reasoning Disentanglement. In order to disentangle perception from reasoning, we add the ground-truth 2D BBox for all objects to the benchmark. This allows repeating our experiments with BBox overlaid, effectively isolating the reasoning and perception. As expected, adding GT BBox as visual prompts to the images will help the VLM to handle perception easier, therefore improving the results. The results in Table 7 show that the major limitation of baselines is not 2D perception, but 3D perception and reasoning. More reasoning analysis on Ego3D-Bench as well as Ego3D-VLM’s robustness in challenging conditions are given in Appendix A.3 and A.6.
Cognitive Map Format. We explore 3 different formats for our generated cognitive map: visual, JSON, and textual. Examples of the visual and textual maps are shown in Figure 5. The average results are presented in Table 8. Among the three, the textual and JSON formats achieve the best overall performance.
6 Conclusion and Future Work
In this work, we proposed Ego3D-Bench, a benchmark for spatial understanding of VLMs on ego-centric multi-view images. Overall, the benchmark shows a significant gap between human scores and VLMs. To address this limitation, we provided a post-training solution named Ego3D-VLM to enhance the performance of VLMs. Future work should explore fine-tuning VLMs with ego-centric multi-view QAs and incorporating 3D projection modules proposed in Ego3D-VLM in the course of fine-tuning. The limitations of this work are provided in the Appendix (A.9).
Appendix A Appendix
In this appendix, we present the following additional discussions and experimental results corresponding to our proposed Ego3D-Bench and Ego3D-VLM:
  • Ego3D-Bench vs. Other Benchmarks
  • Further details on Ego3D-Bench creation and question templates
  • Benchmarking the thinking of Ego3D-Bench as an Open-QA
  • Results of the three source datasets: NuScenes, Waymo, and Argoverse
  • Benchmarking more generalist VLMs
  • Inference time analysis of Ego3D-VLM vs. other baselines
  • Robustness of Ego3D-VLM in challenging conditions
  • Qualitative results and prompts structure
  • Limitations and future work
A.1 Ego3D-Bench vs. Other Benchmarks
As described in the main body of this paper, our proposed Ego3D-VLM is designed for ego-centric multi-view scenarios, which are particularly relevant for AI agent applications. Ego3D-VLM is different from alternative multi-view settings, specifically those used in All-Angle Bench [46] and VSI-Bench [44]. Figure 7 illustrates the configurations of these benchmarks in comparison to Ego3D-Bench. All-Angle Bench features stationary multi-view cameras observing the same scene from different angles—a setup commonly used in motion capture and surveillance systems. VSI-Bench, on the other hand, involves a single moving camera within a static indoor environment, typically used for static scene reconstruction and SU. In contrast, our Ego3D-Bench assumes multiple cameras mounted on a moving ego agent, capturing the surrounding environment—a configuration suited for AI agent applications such as robotics and autonomous driving.
Figure 7: Settings of multi-view/video spatial understanding benchmarks for VLMs.
A.2 Ego3D-Bench: Further Details
In this section, we provide the templates used to create questions of the benchmark. Figure 8 shows templates used to create each category of Ego3D-Bench. We replace <obj1>, <obj2>, and <obj3> in the templates with object descriptions. <view1>, <view2>, and <view3> are replaced with the camera views from which the object is visible (e.g., Front-Right, Left, etc.). The placeholder <direction> in motion reasoning tasks is substituted with one of the four cardinal directions—"north", "east", "west", or "south"—and this process is repeated until all directions have been used. The placeholder <Y> is for motion reasoning tasks and is replaced with a number between 2 to 5 meters. <X1>, <X2>, <X3>, and <X4> serve as placeholders for multiple-choice options in absolute distance estimation tasks. One of these options is replaced with the ground-truth distance, while the remaining are filled with randomly generated values, ensuring that the distance between any two options is at least 8 meters.
Figure 8: Templates used to create questions of Ego3D-Bench.
A.3 Benchmarking the Thinking of Ego3D-Bench as an Open-QA
All our question-answer pairs include a "thinking" step in open-ended format. In this section, we generate "GT thinking" for further open-ended type evaluation. To this end, we used the GT cognitive maps along with GPT-4o to generate the reasoning for two categories that require numerical answers: Ego-Centric / Object-Centric Absolute Distance (in meters). Table 9 shows that Ego3D-VLM thinking is better than the baselines.
Table 9: Benchmarking the thinking of VLMs on Ego3D-Bench as an open-ended category. The GT thinking is generated using GT cognitive maps along with GPT-4o.
Thinking of Obj. Abs. Dist. Thinking of Ego Abs. Dist.
$\uparrow GPT-4o Score (0-10)\uparrow$
InternVL3-8B 1.9 2.0
Ego3D-VLM InternVL3-8B 4.7 5.2
InternVL3-14B 2.3 2.4
Ego3D-VLM InternVL3-14B 5.7 5.0
InternVL3-38B 2.1 2.5
Ego3D-VLM InternVL3-38B 7.1 3.7
InternVL3-78B 2.1 1.7
Ego3D-VLM InternVL3-78B 7.0 6.8
A.4 Benchmarking Generalist VLMs: Further Results
In the main paper, we reported results for InternVL3, Qwen2.5-VL, and Ovis-2 as representative open-source models, as they are, to the best of our knowledge, the current SOTA among open-source VLMs. Table 10 extends this comparison by including additional generalist VLMs on Ego3D-Bench, such as Gemini-1.5-Flash [9], Gemini-2-Flash [9], Gemini-2.5-Flash [9], Phi-3.5 [12], LLaVA-One-Vision-7B [21], LLaVA-Next-Video-7B [48], and DeepSeek-VL2 [43]. We observe that LLaVA-One-Vision, Phi-3.5, and DeepSeek-VL2 underperform compared to InternVL3 and Qwen2.5-VL models of similar size. For instance, InternVL3-8B achieves an accuracy of 43%, while LLaVA-One-Vision-7B achieves 38.7%. Likewise, Phi-3.5 (3.8B) attains an accuracy of 40.3%, compared to 41% from Ovis-2 (4B).
Table 10: Results of further generalist VLMs on Ego3D-Bench.
[…]
Ego
Obj.
Ego
Obj.
Travel
Ego
Obj.
Ego
Obj.
Model
Dist.
Dist.
Loc.
Mot.
Mot.
Time
Rel.
Rel.
[…]
Dist.
Dist.
[…]
Gemini
Gemini-1.5 Flash 40.4 26.9 50.5 70.1 61.8 28.0 53.6 75.6 50.9 10.8 31.0 20.9
Gemini-2-Flash 43.3 38.6 54.1 58.0 39.3 29.1 47.1 68.5 47.2 10.9 29.4 20.1
Gemini-2.5-Flash 41.3 25.6 65.1 93.4 84.8 20.1 63.5 68.5 57.8 11.4 19.6 15.5
Phi
Phi3.5 28.3 30.8 24.2 59.6 55.9 21.3 45.3 56.7 40.3 22.8 49.4 36.1
LLaVA Family
LLaVA-Next-Video-7B 30.2 26.6 42.2 60.8 58.0 48.3 55.1 73.5 49.3 19.5 22.8 21.2
LLaVA-OV-7B 23.5 22.1 58.4 56.3 52.5 17.9 59.0 20.0 38.7 17.8 19.3 18.5
DeepSeek Family
DeepSeek-VL2-tiny 34.7 29.0 19.9 60.0 59.3 46.6 57.8 65.0 46.5 14.0 18.7 16.3
DeepSeek-VL2-Small 18.1 19.2 28.6 50.3 55.6 37.4 50.5 59.3 39.9 12.7 14.7 13.7
DeepSeek-VL2 22.3 25.3 32.4 60.0 59.1 21.4 56.7 62.0 42.4 20.3 17.4 18.8
A.5 Inference Time Analysis
Table 11 shows the inference time analysis of Ego3D-VLM compared to baseline models. We report End-to-End Latency (E2E Lat.) in seconds and Peak Memory in GB. For the E2E Lat., we measure the average end-to-end inference of models on 50 samples of Ego3D-Bench and for the peak memory we report the peak memory usage during inference on the same samples. Experiments are performed using flash-attention-2. The memory and latency overhead of Ego3D-VLM over InternVL3-78B is 0.6% and 31% , respectively. The main reason for the latency overhead is that the model reasons more when cognitive map is provided in the Ego3D-VLM. In order to make the inference more efficient and deal with the latency overhead of the VLMs, post-training techniques such as quantization [15], token pruning [1], or knowledge distillation [13] can be used.
Table 11: Inference time and memory usage of Ego3D-VLM compared to the baselines.
E2E Lat. Memory E2E Lat. Memory
(sec) (GB) (sec) (GB)
InternVL3-8B 5.2 18.1 InternVL3-38B 16.9 80.0
Ego3D-VLMInternVL3-8B 8.6 26.5 Ego3D-VLMInternVL3-38B 19.1 84.6
InternVL3-14B 15.5 33.1 InternVL3-78B 35.0 161.7
Ego3D-VLMInternVL3-14B 16.4 40.2 Ego3D-VLMInternVL3-78B 46.9 162.4
A.6 Robustness of Ego3D-VLM in Challenging Conditions
To isolate tool performance, we included an ablation with ground-truth cognitive maps in the main body of the paper (Table 6), which improve the results from 60.1% to 79.4% accuracy. This confirms that, there is still a gap between the accuracy of such imperfect tools and ground-truth REC/depth. However, compared to the baseline, our solution still achieves more robust results even in challenging conditions such as low brightness, motion blur, and occlusion. To support this claim, we simulated these conditions and re-ran more specific experiments shown in the Table 12. Although G-DINO and Depth-Anything used in Ego3D-VLM are among the best tools for REC and depth estimation, our solution is orthogonal to any such tools, and has no limitation in this regard.
Table 12: Robustness of external tools in scenarios involving occlusion, motion blur, and low light.
Mult. Choice Abs. Ans.
[…]
InternVL3-8B 43.1 27.2
Ego3D-VLM InternVL3-8B 60.1 8.0
InternVL3-8B [60% Darkness] 41.1 29.8
Ego3D-VLM InternVL3-8B [60% Darkness] 59.6 10.6
InternVL3-8B [Motion Blur, 15×1 kernel] 42.5 28.5
Ego3D-VLM InternVL3-8B [motion blur, 15×1 kernel] 57.9 9.9
InternVL3-8B [30% Occlusion] 42.0 28.9
Ego3D-VLM InternVL3-8B [30% Occlusion] 58.7 10.7
A.7 Results of different Source Datasets
Table 13 presents the results for different source dataset splits used to create Ego3D-Bench. Despite the varying number of camera viewpoints across the three datasets, the performance deviations are minimal. This consistency underscores the reliability of Ego3D-Bench as a benchmark, indicating that model performance is not heavily influenced by the specific choice of source data—a desirable property for robust and fair benchmark design.
Table 13: Results on samples of Ego3D-Bench generated from each of the source datasets.
NuScenes Waymo Argoverse
[…]
Closed-source Models
GPT-4o 60.9 21.2 60.0 19.0 59.4 17.1
Ego3D-VLM GPT-4o 73.5 8.5 73.1 7.5 72.7 6.2
Gemini-1.5-Pro 59.3 26.4 55.6 12.7 57.5 19.5
Ego3D-VLM Gemini-1.5-Pro 77.1 8.1 68.9 6.3 73.0 7.1
Qwen2.5 Family
Qwen2.5-3B 39.7 33.4 38.6 31.5 38.9 31.2
Ego3D-VLM Qwen2.5-3B 44.7 13.0 40.3 12.1 48.1 12.6
Qwen2.5-7B 40.9 36.0 40.6 24.7 41.6 30.1
Ego3D-VLM Qwen2.5-7B 56.4 11.3 51.9 8.3 54.4 8.7
Qwen2.5-32B 56.6 17.5 56.0 11.0 59.2 18.7
Ego3D-VLM Qwen2.5-32B 64.3 17.1 64.8 10.6 64.0 12.1
Qwen2.5-72B 58.2 21.2 57.7 11.1 58.2 16.2
Ego3D-VLM Qwen2.5-72B 74.4 8.5 64.4 6.4 69.5 7.6
InternVL3 Family
InternVL3-8B 42.3 28.3 44.0 21.7 42.9 31.5
Ego3D-VLM InternVL3-8B 62.0 10.5 57.1 6.5 61.3 6.7
InternVL3-14B 51.7 20.4 50.2 15.8 53.3 16.6
Ego3D-VLM InternVL3-14B 69.0 8.8 63.7 7.5 67.0 6.3
InternVL3-38B 51.2 31.4 50.5 19.4 53.4 25.4
Ego3D-VLM InternVL3-38B 71.7 8.7 64.3 8.2 67.0 6.3
InternVL3-78B 60.0 16.7 59.7 10.7 60.1 13.7
Ego3D-VLM InternVL3-78B 76.5 8.3 67.0 6.5 71.0 7.5
Ovis2 Family
Ovis2-4B 41.2 29.2 43.0 18.1 40.9 22.5
Ego3D-VLM Ovis2-4B 47.3 9.8 47.5 7.1 49.7 8.3
Ovis2-8B 41.5 21.4 41.9 20.7 41.0 20.2
Ego3D-VLM Ovis2-8B 55.6 9.4 55.1 6.8 57.9 7.0
Ovis2-16B 48.6 16.2 45.5 11.9 46.8 12.9
Ego3D-VLM Ovis2-16B 62.6 8.5 57.9 7.15 62.2 6.5
A.8 Qualitative Results and Prompts Structure
Figures 9-19 demonstrate example responses of InternVL3-78B and Ego3D-VLMInternVL3-78B on all categories of Ego3D-Bench. As seen, Ego3D-Bench enhances the spatial reasoning ability of the baseline by providing the textual cognitive map.
A.9 Limitations and Future Work
This work has several limitations that need further investigation. The proposed Ego3D-VLM relies on the reasoning capabilities of the underlying vision-language models (VLMs). Consequently, for models with limited reasoning ability (e.g., VILA1.5-8B), we observe little to no improvement when combined with Ego3D-VLM.
Additionally, the REC models used in our pipeline provide 2D bounding box locations for all expressions in the prompt. This results in redundant information within the cognitive maps, which may confuse or mislead the VLMs. Our ablation studies demonstrated that when the list of objects mentioned in the prompt is known in advance, the RMSE improves by approximately 1.5 meters.
Another limitation lies in the accuracy of metric depth estimation in outdoor environments. To mitigate this, we proposed a relational scaling approach based on common-sense object sizes. However, this method is inherently approximate and not fully reliable. Future work should aim to address these issues by enhancing the depth understanding of VLMs or by improving metric depth estimation in outdoor settings. Better integration of spatial reasoning and scene geometry into VLMs could further improve performance in complex 3D environments. Another major limitation in spatial reasoning datasets and benchmarks is the availability of ground-truth 3D. Previous work has shown the effectiveness of neural radiance fields in generating photo-realistic 3D data [14]. Future work should investigate this direction for extending spatial reasoning dataset without having ground-truth 3D annotations.
A.10 Statement on LLMs Assistance
We declare that some portions of this document have been lightly refined using Large Language Models (e.g., ChatGPT) to enhance clarity and polish. All substantive content and ideas remain entirely our own.
Figure 9: Example responses of the baseline and Ego3D-VLM on ego-centric absolute distance task.
Figure 10: Example responses of the baseline and Ego3D-VLM on object-centric absolute distance task.
Figure 11: Example responses of the baseline and Ego3D-VLM on multi-choice ego-centric absolute distance task.
Figure 12: Example responses of the baseline and Ego3D-VLM on multi-choice object-centric absolute distance task.
Figure 13: Example responses of the baseline and Ego3D-VLM on object-centric motion reasoning task.
Figure 14: Example responses of the baseline and Ego3D-VLM on ego-centric motion reasoning task.
Figure 15: Example responses of the baseline and Ego3D-VLM on travel time (ego-centric) task.
Figure 16: Example responses of the baseline and Ego3D-VLM on travel time (object-centric) task.
Figure 17: Example responses of the baseline and Ego3D-VLM on object-centric relative distance task.
Figure 18: Example responses of the baseline and Ego3D-VLM on ego-centric relative distance task.
Figure 19: Example responses of the baseline and Ego3D-VLM on localization task.